# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 4. Proof of Structure Theorem for Minimal Initial Functors
For any $d\in D$, the categories $(F\downarrow d)$ and $(\operatorname{thin}(F)\downarrow d)$ have the same objects, since $C$ and $\operatorname{thin}(C)$ have the same objects, while $F$ and $\operatorname{thin}(F)$ act identically on objects. In particular, since $(F\downarrow d)$ is nonempty, so is $(\operatorname{thin}(F)\downarrow d)$. To show that $(\operatorname{thin}(F)\downarrow d)$ is connected, consider $(c,\gamma),(c^{\prime},\gamma^{\prime})$ in $(\operatorname{thin}(F)\downarrow d)$. These are also objects in the connected category $(F\downarrow d)$, so there is a path in $(F\downarrow d)$ connecting $(c,\gamma)$ and $(c^{\prime},\gamma^{\prime})$. Applying $\pi$ yields a path in $(\operatorname{thin}(F)\downarrow d)$ connecting $(c,\gamma)$ and $(c^{\prime},\gamma^{\prime})$. Hence $(\operatorname{thin}(F)\downarrow d)$ is connected. ∎
@@
If $A$ is connected and $x,y\in A$, then there is a path in $A$ of the form
[…]
Let $r_{i}$ denote the $i^{\mathrm{th}}$ relation in this path, starting from the left. By the way $\mathcal{W}$ is defined, $r_{i}$ belongs to some $P_{i}\in\mathcal{W}$. Then the sequence $P_{1},\ldots,P_{l}$ has the required properties. Conversely, assume that for each $x,y\in A$ there exists a sequence of $P_{1},\ldots,P_{l}$ of subposets in $\mathcal{W}$ as in the theorem statement. Choose a sequence of elements $x=a_{0},a_{1},\ldots,a_{l}=y$, such that $a_{i}\in P_{i}\cap P_{i+1}$ for $i\in\{1,\ldots,l-1\}$. Then for each $i\in\{1,\ldots,l\}$, we have $a_{i-1},a_{i}\in P_{i}$ so since $P_{i}$ is connected, there exists a path in $P_{i}$ from $a_{i-1}$ to $a_{i}$. These paths assemble into a path in $A$ from $x$ to $y$. Hence $A$ is connected. ∎
@@
    an initial scaffold $j\colon P\hookrightarrow Q$ exists,
@@
    the inclusion $j$ is a minimal initial functor,
@@
    if $Q$ has a finite initial scaffold, then for any minimal initial functor $F\colon C\to Q$, there is an initial scaffold $j\colon P\hookrightarrow Q$ and a unique isomorphism $E\colon C\to P$ such that $F=j\circ E$.
@@
Since $Q$ has finite downsets, for all $q\in Q$, each component $A$ of $\operatorname{dn}(q)$ contains a minimum. Hence an initial scaffold $j\colon P\to Q$ always exists; this gives (i). Alternatively, (i) follows from Proposition˜4.3 (ii), since there always exists an initial functor with target $Q$, e.g., the identity functor.
In view of Proposition˜2.12, showing that $j$ is initial amounts to checking that for each $q\in Q$, the intersection $\operatorname{dn}[q]\cap P$ is connected. We proceed by induction on $q$ with respect to the partial order on $\operatorname{dn}[q]$; such an induction makes sense because $Q$ has finite downsets. The base case is that $q$ is a minimum. Then $q\in P$, so
@@
be the connected components of $\operatorname{dn}(q)$.
We claim that for each $i$, $A_{i}\cap P$ is connected. Since $Q$ has finite downsets, there is a finite set $S\subseteq A_{i}$ such that
$A_{i}=\bigcup_{s\in S}\operatorname{dn}[s].$
Since $A_{i}$ is connected, as is each downset $\operatorname{dn}[s]$, Lemma˜4.6 implies that for each $x,y\in A_{i}\cap P$ there is a sequence $s_{1},\ldots,s_{l}$ of elements in $S$ such that $x\in\operatorname{dn}[s_{1}]$, $y\in\operatorname{dn}[s_{l}]$, and
$\operatorname{dn}[s_{i}]\cap\operatorname{dn}[s_{i+1}]$
is non-empty for each $i$. Each such intersection contains a minimum of $Q$ and hence an element of $P$, since $P$ contains all minima. Thus, letting $P_{i}=\operatorname{dn}[s_{i}]\cap P$, each intersection $P_{i}\cap P_{i+1}$ is non-empty. By the induction hypothesis, each $P_{i}$ is connected. Noting that
$A_{i}\cap P=\left(\bigcup_{s\in S}\operatorname{dn}[s]\right)\cap P=\bigcup_{s\in S}\operatorname{dn}[s]\cap P,$
applying Lemma˜4.6 with $A=A_{i}\cap P$ and $\mathcal{W}=\{\operatorname{dn}[s]\cap P\}_{s\in S}$ gives that $A_{i}\cap P$ is connected, as claimed.
@@
by inserting $q$, together with the relations $m_{i}<q$ for each $i$. Since each $A_{i}\cap P$ is connected, it follows that $\operatorname{dn}[q]\cap P$ is connected. This finishes the proof that $j$ is initial.
To see that $j$ is minimal, note that for any initial functor $F\colon C\to Q$, Proposition˜4.3 (i) implies that $|\operatorname{Ob}P|=|I_{Q}|\leq|\operatorname{Ob}C|$. Moreover, Proposition˜4.3 (ii) implies that for every $q\in E_{Q}$ and component $A$ of $\operatorname{dn}(q)$ there is a relation $a<q$ in $\operatorname{im}F$ with $a\in A$. It follows that $|\hom P|\leq|\hom C|$. Hence $j$ is minimal. This proves (ii).
To prove (iii), assume that $Q$ has a finite initial scaffold (i.e., $I_{Q}$ is finite) and let $F\colon C\to Q$ be a minimal initial functor. Then (ii) and Proposition˜4.3 (i) together imply that $F$ is an injection on objects. In addition, Lemma˜4.5 implies that $C$ is thin, since if $C$ is not already thin, then $|\hom(\operatorname{thin}(C))|<|\hom C|$, contradicting that $F$ is minimal. Therefore, $F$ is injective on both objects and morphisms, so $\operatorname{im}F$ is a subcategory of $Q$, and in particular, is closed under composition. That is, $\overline{\operatorname{im}F}=\operatorname{im}F$. By Proposition˜4.3 (ii), $\overline{\operatorname{im}F}=\operatorname{im}F$ contains an initial scaffold $P$. The minimality of $F$ now implies that $\operatorname{im}(F)=P$. It follows that we have a unique isomorphism $E\colon C\to P$ satisfying $F=j\circ E$; namely, $E$ is the functor obtained from $F$ by restricting the target to $\operatorname{im}F=P$. ∎
@@
Given a poset $Q$ with finite downsets, an initial scaffold $P\subseteq Q$, and $q\in Q$, the inclusion $P\cap\operatorname{dn}(q)\hookrightarrow\operatorname{dn}(q)$ induces a bijection on path components.
@@
If $q$ is a minimum, then $\operatorname{dn}(q)$ is empty, so this is clear. If $q$ is not a minimum, the statement is equivalent to the statement that for each connected component $A_{i}$ of $\operatorname{dn}(q)$, $P\cap A_{i}$ is connected. We have shown this in the inductive step of the proof that $P\hookrightarrow Q$ is initial (Theorem˜3.5 (ii)). ∎
@@ 5. Size of Initial Scaffolds of Intervals in $\mathbb{N}^{d}$
In this section, we prove Theorem˜3.10. Recall that this theorem gives tight bounds on the size of an initial scaffold of an interval $Q\subseteq\mathbb{\mathbb{N}}^{d}$, in terms of the number of minima of $Q$. The proof uses well-known results about the Betti numbers of monomial ideals. We begin by reviewing these.
@@ 5.1. Betti Numbers of Monomial Ideals
A standard structure theorem [peeva2010graded, Sections 1.7 and 1.9] says that if $G\colon\mathbb{N}^{d}\to\mathbf{Vec}$ is finitely generated, then there exists a minimal resolution $X$ of $G$ which is unique up to isomorphism, and each $X_{i}$ is finitely generated. Moreover, Hilbert’s Syzygy Theorem says that $X_{i}=0$ for $i>d$. Recalling the notation of Section˜2.4.1, we call the function $\beta^{X_{i}}\colon\mathbb{N}^{d}\to\mathbb{N}$ the $i^{\mathrm{th}}$ (multigraded) Betti number of $G$, and denote it as $\beta_{i}^{G}$.
@@
We regard a monomial ideal $J$ as an $\mathbb{N}^{d}$-indexed persistence module with
@@
A version of a well-known result called Hochster’s formula [miller2005combinatorial, Theorem 1.34] gives the Betti numbers of a monomial ideal $J$ in terms of the reduced homology of a simplicial complex built from $J$. To state the result, we need the following notation and definition: Given $s\subseteq\{1,...,d\}$, let $\mathbf{e}_{s}=\sum_{j\in s}\mathbf{e}_{j}\subseteq\{0,1\}^{d}$, where $\mathbf{e}_{j}\subseteq\mathbb{N}^{d}$ denotes the $j^{\mathrm{th}}$ standard basis vector.
@@
For $J\subseteq R^{d}$ an ideal and $z\in\mathbb{N}^{d}$, define the Upper Koszul complex of $J$ at $z$ to be the simplicial complex
@@
Thus, vertices in $K^{z}(J)$ are indices $j\subseteq\{1,\ldots,d\}$ such that $x^{z-\mathbf{e}_{j}}\in J$ and edges are pairs of coordinates $\{j,k\}$ such that $x^{z-\mathbf{e}_{j}-\mathbf{e}_{k}}\in J$. In the following, $\tilde{H}$ denotes reduced simplicial homology with coefficients in $\mathbbm{k}$.
Theorem 5.3 ([miller2005combinatorial, Theorem 1.34]).
@@ 5.2. Bounds on Betti Numbers of Monomial Ideals
A well-known result of Bayer, Peeva, and Sturmfels [bayer1998monomial] provides bounds on the Betti numbers of monomial ideals. See also [miller2005combinatorial] for an expository treatment. We state an asymptotic version of the result, leveraging a standard upper bound on the number of $i$-dimensional faces of a convex polytope [edelsbrunner1987algorithms, Theorem 6.12].
Theorem 5.5 ([bayer1998monomial]).
Suppose $J\subseteq R^{d}$ is a monomial ideal with a minimal generating set of size $n$. Then for each $i$,
@@
Theorem˜5.5 is proven by showing that $\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{i}(z)$ is at most the number of $i$-dimensional faces of a polytope in $\mathbb{R}^{d}$ with $n$ vertices. McMullen’s famous upper bound theorem [mcmullen1970maximum], which asserts that the maximum number of $i$-dimensional faces is attained by a cyclic polytope, gives a tight, non-asymptotic bound on the number of such faces. This leads to a non-asymptotic version of Theorem˜5.5, which appears in [bayer1998monomial, miller2005combinatorial]. By instead using the asymptotic bound on the number of $i$-dimensional faces given in [edelsbrunner1987algorithms, Theorem 6.12], we obtain Theorem˜5.5.
@@
$\sum_{z\in\mathbb{N}^{2}}\beta^{J}_{1}(z)=n-1.$
Thus, the bound $\Theta(n)$ holds for $d=2$. Example˜5.14 below gives a family of monomial ideals in $\mathbb{N}^{4}$ for which the sum of first Betti numbers grows quadratically with the number of minima. Thus, the bound $\Theta(n^{2})$ holds for $d=4$. Finally, for any monomial ideal $J\subseteq R^{d}$ and $d^{\prime}>d$, note that $J^{\prime}\coloneqq J\times R^{d^{\prime}-d}$ is an ideal of $R^{d^{\prime}}$ such that for all $i$,
@@ 5.3. Size Bounds
Our proof of Theorem˜3.10 will make essential use of the following result, which will also be useful in Sections˜6.3 and 6.2, when we consider the problem of computing an initial scaffold of an interval in $\mathbb{N}^{d}$.
All upsets below are taken in the poset $\mathbb{N}^{d}$.
@@
For $Q\subseteq\mathbb{N}^{d}$ an interval, $I_{Q}=I_{\operatorname{up}[Q]}\cap Q$. In particular, $I_{Q}\subseteq I_{\operatorname{up}[Q]}$.
@@
A poset $U\subseteq\mathbb{N}^{d}$ is an upset if $U=\operatorname{up}[S]$ for some set $S\subseteq\mathbb{N}^{d}$. Given an upset $U\subseteq\mathbb{N}^{d}$, define a monomial ideal
$\hat{U}=\langle x^{z}\mid z\in U\rangle\subseteq R^{d}.$
@@
The map $U\mapsto\hat{U}$ is a bijection between the set of upsets in $\mathbb{N}^{d}$ and the set of monomial ideals of $R^{d}$.
@@
For $U\subseteq\mathbb{N}^{d}$ an upset and $z\in U$, let $L$ denote the full subposet of $\operatorname{dn}(z,U)$ with
$\operatorname{Ob}L=\{y\in\operatorname{dn}(z,U)\mid\|y-z\|_{1}\leq 2\}.$
The inclusion $L\hookrightarrow\operatorname{dn}(z,U)$ induces a bijection on path components.
@@
In what follows, we write $D=\operatorname{dn}(z,U)$. Since $U$ is an interval, each component of $D$ contains an element of the form $z-\mathbf{e}_{i}$ for some $\mathbf{e}_{i}$. Since $z-\mathbf{e}_{i}\in L$, the inclusion $L\hookrightarrow D$ induces a surjection on path components. It remains to show that this inclusion is an injection.
Consider $w,y\in L$ lying in the same component of $D$. We need to show that $w,y$ lie in the same component of $L$. There exists a path $w=q_{1},\ldots,q_{n}=y$ from $w$ to $y$ in $D$. We will modify this path to obtain a path from $w$ to $y$ lying in $L$. Clearly, we may assume $n>2$. We may assume without loss of generality that for each triple $q_{i-1},q_{i},q_{i+1}$ of consecutive points in the path, either
$q_{i-1}<q_{i}>q_{i+1}\quad\textup{or}\quad q_{i-1}>q_{i}<q_{i+1}.$
In the former case, we call $q_{i}$ an upper point, and in the latter case, we call $q_{i}$ a lower point.
If $w=z-\mathbf{e}_{j}-\mathbf{e}_{j^{\prime}}$ (for some $j$ and $j^{\prime}$), then by inserting a new element $w-\mathbf{e}_{j}$ in the path, if necessary, we may assume without loss of generality that $w<q_{2}$. Symmetrically, if $y=z-\mathbf{e}_{j}-\mathbf{e}_{j^{\prime}}$, then we may assume that $q_{n-1}>y$.
Since $U$ is an interval, we may further assume that each upper point $q_{i}$ is of the form $z-\mathbf{e}_{j}$, by replacing $q_{i}$ with a point above it, if necessary. Given these assumptions, for each lower point $q_{i}$, we have $q_{i-1}=z-\mathbf{e}_{j}$ and $q_{i+1}=z-\mathbf{e}_{j^{\prime}}$, so we have
@@
where $\wedge$ denotes the meet. It follows that $\wedge(q_{i-1},q_{i+1})\in L$. Therefore, by replacing each lower point $q_{i}$ with $\wedge(q_{i-1},q_{i+1})$, we obtain a path from $w$ to $y$ in $L$. This shows that $w,y$ lie in the same component of $L$. Hence the map on components induced by the inclusion $L\hookrightarrow D$ is an injection, as desired. ∎
@@
By Corollary˜5.4, it suffices to check that $z\in E_{U}$ if and only if $|\pi_{0}(K^{z}(U))|\geq 2$. Note that $z\in E_{U}$ if and only if $\operatorname{dn}(z,U)$ has at least two components. Therefore, it suffices to exhibit a bijection between the components of $\operatorname{dn}(z,U)$ and those of $K^{z}(U)$. Lemma˜5.11 provides a bijection between the components of $\operatorname{dn}(z,U)$ and those of the subposet $L$ in the statement of that lemma. Moreover, components of $L$ correspond bijectively with components of the undirected graph $G$ underlying the Hasse diagram of $L$. Let $K^{+}$ denote the barycentric subdivision of the 1-skeleton of $K^{z}(U)$. The simplicial map $f\colon G\to K^{+}$ given by $f(z-\mathbf{e}_{s})=s$ is a graph isomorphism, and therefore induces a bijection on components. Composing these bijections yields a bijection between the components of $\operatorname{dn}(z,U)$ and those of $K^{z}(U)$. ∎
@@
We identify a family of upsets $(U^{k})_{k\in\mathbb{N}}$ in $\mathbb{N}^{4}$ such that $U^{k}$ has $\Theta(k)$ minima and $|I_{U^{k}}|=\Theta(k^{2})$. By Lemma˜5.13, this corresponds to a family $(J^{k})_{k\in\mathbb{N}}$ of ideals of $R^{4}$ for which
$\sum_{z\in\mathbb{N}^{4}}\beta^{J^{k}}_{1}(z)=\Theta(|M_{J^{k}}|^{2}),$
as we needed in the proof of Corollary˜5.7.

### v2

@@ 4. Proof of Structure Theorem for Minimal Initial Functors
For any $d\in D$, the categories $(F\downarrow d)$ and $(\operatorname{thin}(F)\downarrow d)$ have the same objects because $C$ and $\operatorname{thin}(C)$ have the same objects, while $F$ and $\operatorname{thin}(F)$ act identically on objects. In particular, since $(F\downarrow d)$ is nonempty, so is $(\operatorname{thin}(F)\downarrow d)$. To show that $(\operatorname{thin}(F)\downarrow d)$ is connected, consider $(c,\gamma),(c^{\prime},\gamma^{\prime})$ in $(\operatorname{thin}(F)\downarrow d)$. These are also objects in the connected category $(F\downarrow d)$, so there is a path in $(F\downarrow d)$ connecting $(c,\gamma)$ and $(c^{\prime},\gamma^{\prime})$. Applying $\pi$ yields a path in $(\operatorname{thin}(F)\downarrow d)$ connecting $(c,\gamma)$ and $(c^{\prime},\gamma^{\prime})$. Hence $(\operatorname{thin}(F)\downarrow d)$ is connected. ∎
@@
If $A$ is connected and $x,y\in A$, then there is a path
$x=a_{1},\ldots,a_{l+1}=y$
in $A$. By definition, $a_{i}$ and $a_{i+1}$ are comparable for each $i\in\{1,\ldots,l\}$. Since $A=\bigcup_{P\in\mathcal{W}}P$, each pair $\{a_{i},a_{i+1}\}$ belongs to some $P_{i}\in\mathcal{W}$. The sequence $P_{1},\ldots,P_{l}$ then has the desired properties.
Conversely, assume that for each $x,y\in A$ there exists a sequence $P_{1},\ldots,P_{l}$ of subposets in $\mathcal{W}$ as in the theorem statement. Choose a sequence $x=a_{1},a_{2},\ldots,a_{l+1}=y$ such that $a_{i+1}\in P_{i}\cap P_{i+1}$ for all $i\in\{1,\ldots,l-1\}$. Then for each $i\in\{1,\ldots,l\}$, we have $a_{i},a_{i+1}\in P_{i}$ so since $P_{i}$ is connected, there exists a path in $P_{i}$ from $a_{i}$ to $a_{i+1}$. These paths assemble into a path in $A$ from $x$ to $y$. Hence $A$ is connected. ∎
@@
    an initial scaffold $P\subseteq Q$ exists,
@@
    the inclusion $P\hookrightarrow Q$ is a minimal initial functor,
@@
    if $Q$ has a finite initial scaffold and $F\colon C\to Q$ is a minimal initial functor, then $F$ is an embedding and $\operatorname{im}F$ is an initial scaffold of $Q$.
@@
Since $Q$ has finite downsets, for all $q\in Q$, each component $A$ of $\operatorname{dn}(q)$ contains a minimum. Hence an initial scaffold $P\subseteq Q$ always exists; this gives (i). Alternatively, (i) follows from Proposition˜4.4 (ii), since there always exists an initial functor with target $Q$, namely, the identity functor.
In view of Proposition˜2.16, showing that the inclusion $P\hookrightarrow Q$ is initial amounts to checking that for each $q\in Q$, the intersection $\operatorname{dn}[q]\cap P$ is connected. We proceed by induction on $q$ with respect to the partial order on $\operatorname{dn}[q]$; such an induction makes sense because $Q$ has finite downsets. The base case is that $q$ is a minimum. Then $q\in P$, so
@@
be the components of $\operatorname{dn}(q)$.
We claim that $A_{i}\cap P$ is connected for each $i$. To show this, note that $A_{i}$ is connected, as is each downset $\operatorname{dn}[a]$ for $a\in A_{i}$. Thus, applying Lemma˜4.7 with $A=A_{i}$ and $\mathcal{W}=\{\operatorname{dn}[a]\}_{a\in A_{i}}$ gives that for each $x,y\in A_{i}\cap P$ there is a sequence $a_{1},\ldots,a_{l}$ of elements in $A_{i}$ such that $x\in\operatorname{dn}[a_{1}]$, $y\in\operatorname{dn}[a_{l}]$, and
$\operatorname{dn}[a_{i}]\cap\operatorname{dn}[a_{i+1}]$
is non-empty for each $i$. Each such intersection contains a minimum of $Q$ and hence an element of $P$, since $P$ contains all minima. Thus, letting $P_{i}=\operatorname{dn}[a_{i}]\cap P$, each intersection $P_{i}\cap P_{i+1}$ is non-empty. By the induction hypothesis, each $P_{i}$ is connected. Noting that
$A_{i}\cap P=\left(\bigcup_{a\in A_{i}}\operatorname{dn}[a]\right)\cap P=\bigcup_{a\in A_{i}}\operatorname{dn}[a]\cap P,$
applying Lemma˜4.7 with $A=A_{i}\cap P$ and $\mathcal{W}=\{\operatorname{dn}[a]\cap P\}_{a\in A_{i}}$ gives that $A_{i}\cap P$ is connected, as claimed.
@@
by inserting $q$, together with the relations $m_{i}<q$ for each $i$. Since each $A_{i}\cap P$ is connected, it follows that $\operatorname{dn}[q]\cap P$ is connected. This finishes the proof that the inclusion $P\hookrightarrow Q$ is initial.
To see that $P\hookrightarrow Q$ is minimal, note that for any initial functor $F\colon C\to Q$, Proposition˜4.4 (i) implies that $|\operatorname{Ob}P|=|I_{Q}|\leq|\operatorname{Ob}C|$. Moreover, Proposition˜4.4 (ii) implies that for every $q\in E_{Q}$ and component $A$ of $\operatorname{dn}(q)$ there is a relation $a<q$ in $\operatorname{im}F$ with $a\in A$. It follows that $|\hom P|\leq|\hom C|$. This proves (ii).
To prove (iii), assume that $Q$ has a finite initial scaffold (i.e., $I_{Q}$ is finite) and let $F\colon C\to Q$ be a minimal initial functor. Then (ii) and Proposition˜4.4 (i) together imply that $F$ is an injection on objects. In addition, Lemma˜4.6 implies that $C$ is thin, since otherwise we have $|\hom(\operatorname{thin}(C))|<|\hom C|$, contradicting that $F$ is minimal. Therefore, $F$ is faithful, hence an embedding. Thus, we have $\overline{\operatorname{im}F}=\operatorname{im}F$ by Lemma˜4.2. By Proposition˜4.4 (ii), $\overline{\operatorname{im}F}=\operatorname{im}F$ contains an initial scaffold $P$. The minimality of $F$ implies that $\operatorname{im}F=P$. ∎
@@
Given a poset $Q$ with finite downsets, an initial scaffold $P\subseteq Q$, and $q\in Q$, the inclusion $P\cap\operatorname{dn}(q)\hookrightarrow\operatorname{dn}(q)$ induces a bijection on sets of components.
@@
Since $Q$ has finite downsets and $P$ contains all minima of $Q$, each component of $\operatorname{dn}(q)$ has non-empty intersection with $P$. Thus, the map on components induced by the inclusion $P\cap\operatorname{dn}(q)\hookrightarrow\operatorname{dn}(q)$ is a surjection. It is an injection if and only $P\cap A_{i}$ is connected for each component $A_{i}$ of $\operatorname{dn}(q)$. We have shown this in the inductive step of the proof of Theorem˜3.5 (ii). ∎
@@ 5. Size of Initial Scaffolds of Intervals in $\mathbb{N}^{d}$
In this section, we prove Theorem˜3.10. Recall that for an interval $Q\subseteq\mathbb{\mathbb{N}}^{d}$, this theorem gives tight bounds on the size of an initial scaffold of $Q$ in terms of the number of minima of $Q$. The proof uses well-known results about the Betti numbers of monomial ideals. We begin by reviewing these.
@@ 5.1. Betti Numbers of Monomial Ideals
A standard structure theorem [38, Sections 1.7 and 1.9] says that if $G\colon\mathbb{N}^{d}\to\mathbf{Vec}$ is finitely generated, then there exists a minimal resolution $X$ of $G$ that is unique up to isomorphism, with each $X_{i}$ finitely generated. Moreover, Hilbert’s Syzygy Theorem says that $X_{i}=0$ for $i>d$. Recalling the notation of Section˜2.4.1, we denote the function $\beta^{X_{i}}\colon\mathbb{N}^{d}\to\mathbb{N}$ as $\beta_{i}^{G}$ and call its values the $i^{\mathrm{th}}$ (multigraded) Betti numbers of $G$.
@@
We regard a monomial ideal $J$ as an $\mathbb{N}^{d}$-module with
@@
A version of a well-known result called Hochster’s formula [35, Theorem 1.34] gives the Betti numbers of $J$ in terms of the reduced homology of a simplicial complex built from $J$. To state the result, we need the following notation and definition: Given $s\subseteq\{1,...,d\}$, let $\mathbf{e}_{s}=\sum_{j\in s}\mathbf{e}_{j}\in\{0,1\}^{d}$, where $\mathbf{e}_{j}\in\mathbb{N}^{d}$ denotes the $j^{\mathrm{th}}$ standard basis vector.
@@
For $J\subseteq R^{d}$ a monomial ideal and $z\in\mathbb{N}^{d}$, define the Upper Koszul complex of $J$ at $z$ to be the simplicial complex
@@
Thus vertices in $K^{z}(J)$ are indices $j\in\{1,\ldots,d\}$ such that $x^{z-\mathbf{e}_{j}}\in J$ and edges are pairs of indices $\{j,k\}$ such that $x^{z-\mathbf{e}_{j}-\mathbf{e}_{k}}\in J$. In the following, $\tilde{H}$ denotes reduced simplicial homology with coefficients in $\mathbbm{k}$.
Theorem 5.3 ([35, Theorem 1.34]).
@@
It follows from Theorem˜5.3 that $\beta_{i}^{J}=0$ for all $i\geq d$; this can also be seen by applying Hilbert’s syzygy theorem to the quotient module $R^{d}/J$.
@@ 5.2. Bounds on Betti Numbers of Monomial Ideals
A well-known result of Bayer, Peeva, and Sturmfels [5, Theorem 6.3] provides bounds on the Betti numbers of monomial ideals. See also [35] for an expository treatment. We state an asymptotic version of the result, leveraging a standard upper bound on the number of $i$-dimensional faces of a convex polytope [21, Theorem 6.12]. In what follows, we assume $d$ is constant.
Theorem 5.5 ([5]).
If $J\subseteq R^{d}$ is a monomial ideal with a minimal generating set of size $n$, then for each $i$,
@@
Theorem˜5.5 is proven by showing that $\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{i}(z)$ is at most the number of $i$-dimensional faces of a polytope in $\mathbb{R}^{d}$ with $n$ vertices. McMullen’s famous upper bound theorem [34], which states that the maximum number of $i$-dimensional faces is attained by a cyclic polytope, gives a tight, non-asymptotic bound on the number of such faces. This leads to a non-asymptotic variant of Theorem˜5.5, which is what appears in [5]. Instead, using the asymptotic bound on the number of $i$-dimensional faces given in [21, Theorem 6.12], we obtain Theorem˜5.5.
@@
$\sum_{z\in\mathbb{N}^{2}}\beta^{J}_{1}(z)=n-1;$
see [35, Proposition 3.1] for a proof. Thus, the bound $\Theta(n)$ holds for $d=2$. Example˜5.14 below gives a family of monomial ideals in $\mathbb{N}^{4}$ for which the sum of first Betti numbers grows quadratically with the number of minima. Thus, the bound $\Theta(n^{2})$ holds for $d=4$. Finally, for any monomial ideal $J\subseteq R^{d}$ and $d^{\prime}>d$, note that $J^{\prime}\coloneqq J\times R^{d^{\prime}-d}$ is an ideal of $R^{d^{\prime}}$ such that for all $i$,
@@ 5.3. Size Bounds
Our proof of Theorem˜3.10 will make essential use of the following result, which we will also use in Sections˜6.3 and 6.2, where we consider the problem of computing an initial scaffold of an interval in $\mathbb{N}^{d}$.
For what follows, recall the definitions of the sets $I_{Q}$ and $E_{Q}$ given in Sections˜3.1 and 4, respectively. All upsets below are taken in the poset $\mathbb{N}^{d}$.
@@
For any interval $Q\subseteq\mathbb{N}^{d}$, we have $I_{Q}=I_{\operatorname{up}[Q]}\cap Q$. In particular, $I_{Q}\subseteq I_{\operatorname{up}[Q]}$.
@@
A poset $U\subseteq\mathbb{N}^{d}$ is an upset if $U=\operatorname{up}[S]$ for some set $S\subseteq\mathbb{N}^{d}$. Given an upset $U\subseteq\mathbb{N}^{d}$, define the monomial ideal
$\hat{U}\coloneqq\langle x^{z}\mid z\in U\rangle\subseteq R^{d}.$
@@
The map $U\mapsto\hat{U}$ is a bijection between the set of upsets of $\mathbb{N}^{d}$ and the set of monomial ideals of $R^{d}$.
@@
For $U\subseteq\mathbb{N}^{d}$ an upset and $z\in U$, let $T$ denote the full subposet of $\operatorname{dn}(z,U)$ with
$\operatorname{Ob}T=\{y\in\operatorname{dn}(z,U)\mid\|y-z\|_{1}\leq 2\}.$
The inclusion $T\hookrightarrow\operatorname{dn}(z,U)$ induces a bijection on sets of components.
@@
In what follows, we write $D=\operatorname{dn}(z,U)$. Since $U$ is an upset, each component of $D$ contains an element of the form $z-\mathbf{e}_{i}$ for some $\mathbf{e}_{i}$. Since $z-\mathbf{e}_{i}\in T$, the inclusion $T\hookrightarrow D$ induces a surjection on components. It remains to show that this inclusion is an injection.
Consider $w,y\in T$ lying in the same component of $D$. We need to show that $w,y$ lie in the same component of $T$. There exists a path
$w=q_{1},\ldots,q_{k}=y$
from $w$ to $y$ in $D$. By inserting copies of points into the path, we may assume that for each triple $q_{i-1},q_{i},q_{i+1}$ of consecutive points in the path, either
$q_{i-1}\leq q_{i}\geq q_{i+1}\quad\textup{or}\quad q_{i-1}\geq q_{i}\leq q_{i+1}.$
We call $q_{i}$ an upper point in the former case and a lower point in the latter case. Moreover, we may assume that $k>2$ and that $q_{2}$ and $q_{k-1}$ are upper points. Since $U$ is an upset, by replacing each upper point $q_{i}$ with a point above it, if necessary, we may take each upper point to be of the form $z-\mathbf{e}_{j}$. Then for each lower point $q_{i}$, we have $q_{i-1}=z-\mathbf{e}_{j}$ and $q_{i+1}=z-\mathbf{e}_{j^{\prime}}$, so
@@
where $\wedge$ denotes the meet operator. Since $U$ is an upset, it follows that $\wedge(q_{i-1},q_{i+1})\in U$, hence $\wedge(q_{i-1},q_{i+1})\in T$. Therefore, by replacing each lower point $q_{i}$ with $\wedge(q_{i-1},q_{i+1})$, we obtain a path from $w$ to $y$ in $T$. This shows that $w,y$ lie in the same component of $T$, completing the proof. ∎
In what follows, the support of an $\mathbb{N}$-valued function $f$ is denoted $\operatorname{Supp}f$.
@@
By Corollary˜5.4, it suffices to check that $z\in E_{U}$ if and only if $|\pi_{0}(K^{z}(U))|\geq 2$. Note that $z\in E_{U}$ if and only if $\operatorname{dn}(z,U)$ has at least two components. Therefore, it suffices to exhibit a bijection between the components of $\operatorname{dn}(z,U)$ and those of $K^{z}(U)$. Lemma˜5.11 provides a bijection between the components of $\operatorname{dn}(z,U)$ and those of the subposet $T$ in the statement of that lemma. Moreover, components of $T$ correspond bijectively to components of the undirected graph $G$ underlying the Hasse diagram of $T$. Let $K^{+}$ denote the barycentric subdivision of the 1-skeleton of $K^{z}(U)$. The simplicial map $f\colon G\to K^{+}$ given by $f(z-\mathbf{e}_{s})=s$ is a graph isomorphism, and therefore induces a bijection on components. Composing these bijections yields a bijection between the components of $\operatorname{dn}(z,U)$ and those of $K^{z}(U)$. ∎
@@
We identify a family of upsets $(U^{k})_{k\in\mathbb{N}}$ in $\mathbb{N}^{4}$ such that $U^{k}$ has $\Theta(k)$ minima and $\Theta(k^{2})$ essential points. By Lemma˜5.12, this corresponds to a family of ideals in $R^{4}$ for which the sum of first Betti numbers grows quadratically with the number of minima, as needed in the proof of Corollary˜5.7.